a pharmacist wishes to mix a solution that is 2% Minoxidil. She has on hand 70 ml of a 1% solution and wishes to add some 4% solution to obtain the desired 2% solution. how much 4% solution should she add?

Answer by mananth(16946)

--- percent ---------------- Amount
Minoxidil I 1 ---------------- 70 ml
Minoxidil II 4 ---------------- x ml
Mixture 2 ---------------- 70 + x ml

1 * 70 + 4 x = 2 ( 70 + x )
70 + 4 x = 140 + 2 x
4 x -2 x = -70 + 140
2 x = 70
/ 2
x = 35 ml Minoxidil II

m.ananth@hotmail.ca